# Long Grounded Thoughts: Distilling Compositional Visual Reasoning Chains at Scale

### v1

@@ 3 Offline and Online Synthesis for RL
Here, the superscript ++ denotes a correct prediction and - an incorrect one, while $\succ$ indicates preference and $\oplus$ denotes concatenation. Following LPT (liao2025longperceptualthoughts), we collect both positive and negative CoTs from the initial VLM response (e.g., $({\textnormal{z}}_{1}^{+},{\textnormal{a}}_{1}^{+})$ and $({\textnormal{z}}_{1}^{-},{\textnormal{a}}_{1}^{-})$), together with subsequent reasoning expansions from $\mathcal{M}_{\text{Reason}}$ (e.g., $({\textnormal{z}}_{2}^{+},{\textnormal{a}}_{2}^{+})$). Note that we do not explicitly verify whether ${\textnormal{z}}_{1}$ itself is correct or incorrect; instead, we infer its correctness from the associated ${\textnormal{a}}_{1}$, which we find to be a reasonable approximation at scale. For online RL (GRPO), we reuse the same instruction prompts but discard preference responses.
Table 2: Main results on vision-centric reasoning benchmarks. We compare our models against both open- and closed-source VLMs across five challenging benchmarks. Community baselines trained on open source data sometimes underperform the base Qwen2.5-VL-7B-Instruct, underscoring the lack of high-quality open reasoning data for vision-centric tasks. Our models achieve the top score in 3/5 benchmarks (V* Bench, CV Bench, and MMStar-V). Closed source performance is only reported for reference and it was taken from their respective technical reports.
@@
Figure 5: Quantitative and qualitative comparison of the post-training pipeline on our data vs pure RL on the base model. (Right) The graph illustrates the effect of scaling dataset size during online RL. The baseline (blue line), starting from an off-the-shelf model, exhibits negative scaling: performance peaks at 0.695 (10K samples) and degrades with more data. In contrast, our method (green line), which includes SFT on our high-quality data with complex reasoning traces, allows to scale online RL further. This suggests that without offline "skill teaching" via SFT, online RL fails to effectively utilize larger datasets. (Left) A qualitative example (from V* bench), using each model’s best checkpoint (indicated by a dot on the curve), highlights the resulting difference in reasoning. The baseline model fails to identify the partially obscured dog and answers incorrectly. Our model also initially expresses confusion but then self-corrects ("Wait, I’m getting conflicting information…"), showcasing a multi-step reasoning process to arrive at the correct answer. This self-correction capability, instilled with our data, is not observed in the baseline, indicating RL alone was insufficient to elicit this behavior. Image brightness was increased for illustration purposes.
LPT BaseModel 750K None None 0.682
Ours BaseModel 750K None None 0.716
LPT BaseModel 0 GRPO 10K 0.695
50K 0.683
100K 0.669
LPT SFT 750K GRPO 10K 0.662
50K 0.685
100K 0.709
Ours BaseModel 0 GRPO 70K 0.704
SFT 750K DPO 70K 0.737
SFT 750K DPO 129K 0.740
SFT 750K GRPO 70K 0.757
SFT 750K GRPO 129K 0.746
@@ 4 Experiments
In this section, we first evaluate the performance of finetuning a 7B VLM on our data, and compare it compare against three categories: (i) open-weight, open-data models, (ii) open-weight, closed-data models, and (iii) fully closed models. Next, we use our data to systematically analyze the difference post-training stages in a base instruction model, revealing consistent findings. Finally, we evaluate whether our vision-centric reasoning data transfers to a different modality and architecture (e.g. text-only reasoning, embodied QA, and audio reasoning on an Omni-7B model).
4.1 Comparison vs baselines
@@
Base Model and Benchmarks We fine-tune Qwen2.5-VL-7B-Instruct (bai2025qwen25vltechnicalreport), referred to as BaseModel and evaluate on vision-centric tasks. Following liao2025longperceptualthoughts, for general-knowledge datasets, we retain only their vision splits (e.g., MMStar-V (chen2024we)). In addition, we adopt vision-centric focused benchmarks V Bench (v_star), CV-Bench (averaged both 2D and 3D), and MMVP (tong2024cambrian). These benchmarks test visual search, 2D/3D spatial reasoning, fine-grained attribution, and coarse scene understanding. In our table 2, we also use RealWorldQA (RealWorldQA_xAI_2024).
@@
Evaluation. We use two main protocols for evaluation. For comparison vs existing models and published results we used VLMEvalKit (duan2024vlmevalkit) with GPT-4o as a judge (Table 2). For ablations, we utilize the evaluation protocol of  (liao2025longperceptualthoughts) which utilizes rule-based matching instead of a judge. For clarity, results are shown with decimal when no LLM as a judge is used. To avoid ambiguity, results are shown in a 0-1 scale when rule-based evaluation is applied.
Source of high-quality image captions and metadata. We use DOCCI (OnoeDocci2024), a human-annotated dense caption dataset providing comprehensive descriptions of images, including fine-grained visual details. Unlike liao2025longperceptualthoughts, which used only a small subset, we leverage the entire dataset. In addition, each image is processed with the Grounded-Segment-Anything model (ren2024grounded) for additional metadata. We emphasize that our synthesis framework is not tied to DOCCI or any specific object-detection metadata, however exploring other caption datasets is left for future work.
@@
Baselines. We compare against three categories: (i) open-weight, open-data models, (ii) open-weight, closed-data models, and (iii) fully closed models. For open-weight, open-data, we include VLAA-Thinker (chen2025sft), a model trained on $\sim$152K synthetic reasoning traces distilled from a strong reasoning LLM and rewritten with a weaker LLM, using both SFT and RL. We also compare with LPT (liao2025longperceptualthoughts). We also compare vs ReVisual-R1 (chen2025advancingmultimodalreasoningoptimized). Additionally, we include Virgo (du2025virgo), trained on $\sim$19K math-heavy datapoints. For open-weight, closed-data, we consider MiMo-VL-7B, a SoTA VLM trained with a four-stage recipe totaling $\sim$2.4T tokens across images, videos, and text. For fully closed models, we report results for GPT-4o and Claude 3.7 taken from (xiaomi2025mimo).
@@
Training Algorithms. Our posttraining pipeline evaluates both simple stage and multistage SFT as well offline and online RL. SFT. We perform SFT on the large pool of data from Stage 1 and Stage 2. To simplify training, we adopt a curriculum: first fine-tuning on simpler Stage 1 data for one epoch, then introducing the more complex Stage 2 data. Stage 1 consists of $\sim$750K datapoints, and Stage 2 adds $\sim$250K. We fine-tune the language decoder with a batch size of $256$, learning rate $8\times 10^{-6}$. For Stage 1, we train for up to one epoch with maximum image resolution $512\times 512$ and input cutoff length $1024$. For Stage 2, we halve the learning rate and apply early stopping based on validation loss. All experiments are implemented with llama-factory. Offline RL (DPO). We fine-tune the language decoder with a batch size of $256$ using a learning rates of $\{6\times 10^{-6}\}$. Training runs for up to six epochs with early stopping based on validation loss. For DPO, we set $\beta=1$ and, following pang2024iterative, include the SFT loss with a weight of $0.5$. Implemented with llama-factory. Online RL (GRPO). We perform GRPO on top of the fine-tuned model via VERL (verl). We set the batch size 128, max response length 8192 and the learning rate $1\times 10^{-6}$. We use the KL loss coefficient of $0.001$ and omit entropy penalty. To avoid overfitting to a specific answer format, our labels are set in a diverse format (e.g., (A), A, and (A) answer); thus, we compute the reward based on an LLM judge, assigning reward = $1$ to a correct answer and reward = 0 to a wrong answer (as identified by the judge). We additionally assign format reward, we add $0.1$ to the reward if the response is formatted correct (i.e., the model correctly generates both <think> and </think>).
@@
Comparison vs. Baselines. Table 2 compares Qwen2.5-VL-7B fine-tuned with our data against community baselines, close data, and closed models. Models trained on limited open reasoning data (e.g., VLAA-Thinker, Revisual) often underperform even the base Qwen2.5-VL-7B-Instruct, underscoring the lack of high-quality open data for vision-centric reasoning. In contrast, our models consistently improve the base, achieving the best results among open-data approaches and reaching the top score on 3/5 benchmarks.
@@ 4.2 Analysis of VLM Post-training Stages
We analyze VLM post-training strategies using our data. Table 3 reports aggregated performance (average over four benchmarks). Figure 2 shows scaling compared to LPT. For evaluation, we use rule-based matching instead of LLM-as-judge for cost efficiency. Here, we summarize our findings:
Online RL requires prior basic-skill teaching; starting from a base instruct model underperforms SFT on our data. When starting from an off-the-shelf model lacking non-linear thinking patterns (see Figures 54(b)), GRPO peaks at 0.695 (10K) and declines with more data (50–100K), remaining below our SFT baseline (0.716). Finetuning first on LPT data improves over off-the-shelf models but still underperforms simple SFT finetuning on our data. Finetuning on our data leads to the best results suggesting that without offline “skill teaching,” online RL cannot elicit competitive performance vs SFT in high-quality and diverse data. Notably, we do not observed online RL was able to elicit structured complex reasoning traces unless it was pretrained first on our reasoning SFT data (see Figure 5 for a qualitative example).
@@
SFT data diversity matters; grounded MCQ SFT >> LPT SFT. Our grounded-MCQ SFT reaches 0.716 versus 0.682 for LPT SFT (both at comparable scale), a gain of +3.4 points on average. High-diversity, grounded supervision is a strong foundation before either preference tuning or RL. Figure 2 also illustrates a much better scaling when using grounded MCQ SFT data.
@@ 4.3 Beyond Vision-Centric Reasoning
Table 4: Out-of-domain evaluation We evaluate models based on Qwen2.5-VL on out-of-domain tasks, MMLU-Pro (left), and an embodied question answering benchmark (right).
Model Accuracy
Qwen2.5-VL 47.15
VLAA-thinking 21.56
@@
LPT 50.77
Ours (SFT) 50.13
Ours (SFT+DPO) 47.07
Ours (SFT+GRPO) 47.00
Model Accuracy
Qwen2.5-VL 47.55
+ VLAA-thinking 47.85
+ Virgo
+ LPT 51.95
+ Ours (SFT) 48.24
+ Ours (SFT+GRPO) 39.10
+ Ours (SFT+DPO) 56.34
Table 5: Omni-modality out-of-distribution results on audio MMAU Benchmark(Sound, Music, Speech) and text-only (MMLU-Pro). We use Qwen2.5-Omni-7B as baseline Omni model.
@@
In this section, we test whether our vision-centric reasoning data transfers across modalities. Specifically, we evaluate our finetuned model on the text-only benchmark MMLU-PRO (wang2024mmlupro). Table 4 highlights that finetuning in our data does not harm the text-only reasoning capabilities of the model, in fact, we observe notable improvements vs the base model and existing baselines. Despite no containing videos or any embodied-QA data, we also evaluate the model finetuned on our data on a modified version of the NiEH single-evidence QA task (kim2025needlesembodiedhaystackenvironment), observing a notable gains (+10 points) vs base model (Table 4, appendix A.5.3). Finally, we use our data to finetune an Omni Model and evaluate on both the the audio benchmark MMAU (sakshi2024mmau) and MMLU-PRO. As shown in Table 5, fine-tuning with our data produces consistent cross-modal improvements. Relative to the Qwen2.5-Omni-7B baseline, our model (SFT + DPO) achieves gains of +0.98 on MMAU-Sound, +3.02 on MMAU-Music, +0.33 on MMAU-Speech, and +4.07 on MMLU-Pro. These results indicate robust positive transfer across both audio and text modalities, contrasting with Virgo’s negative transfer and LPT’s marginal gains. We attribute this cross-domain generalization to the complex reasoning structures present in our data, which likely promote more generalizable internal representations while preserving alignment with the model’s output distribution. Further experimental details can be found in the suppl material.
@@ 5 Related Work
Recently, several efforts have been made to enhance visual reasoning in VLMs, spanning from test-time techniques (liao-etal-2024-reasoning; liao2024feedbackenhancesemanticgrounding; acuna2025socratic) to improved architectural designs and training pipelines (cheng2024spatialrgpt; wu2025spatial; chen2025eagle25). However, most efforts to distill high-quality reasoning datasets at scale with complex, non-linear behaviors remain largely in the text-only domain and focus primarily on math and code (openthoughts; muennighoff2025s1simpletesttimescaling; jung2025prismatic). In contrast, efforts to build multimodal CoT datasets are growing, yet most reasoning traces remain largely linear and/or utilize existing question rather than synthesizing new ones. For example, LLaVA-CoT (xullavacot) defines reasoning stages in a pipeline—summary, caption, reasoning, and conclusion. SCI-Reason (ma2025scireason) focuses on academic areas and takes a different path, using Monte Carlo Tree Search (MCTS) to collect alternative CoTs. In vision-centric domains such as driving, DrivingVQA (drivingvqa2025) relies on human-annotated traces later enriched with region-level references via GPT-4o, while DriveLMM-o1 (ishaq2025drivelmm) generates CoTs directly by prompting GPT-4o with paired questions and answers. Beyond these linear designs, a smaller body of work studies more complex reasoning structures such as reflection, self-correction, or iterative refinement. For example, Virgo (du2025virgo) distills multimodal CoTs from advanced multimodal reasoning models (qvq2024), focusing on math-heavy domains. VLAA-Thinking (chen2025sft) similarly distills CoTs in the general domain, though subsequent supervised fine-tuning on these traces has been shown to degrade performance. LPT (liao2025longperceptualthoughts) introduces long-form CoTs distilled from large reasoning models, such as R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), with explicit emphasis on expanding thoughts and encouraging models to deliberate more thoroughly across perception-heavy tasks. Our work builds on this line: rather than restricting CoTs to fixed stages or linear narratives or reusing existing MCQs, we synthesize diverse questions and complex reasoning structures, explicitly scaling along three axes: scale, complexity, and reasoning depth.

### v2

@@ 2.4 Offline and Online Synthesis for RL
Here, the superscript ++ denotes a correct prediction and - an incorrect one, while $\succ$ indicates preference and $\oplus$ denotes concatenation. Following previous works, We specifically collect both positive and negative CoTs from the initial VLM response (e.g., $({\textnormal{z}}_{1}^{+},{\textnormal{a}}_{1}^{+})$ and $({\textnormal{z}}_{1}^{-},{\textnormal{a}}_{1}^{-})$), together with subsequent reasoning expansions from $\mathcal{M}_{\text{Reason}}$ (e.g., $({\textnormal{z}}_{2}^{+},{\textnormal{a}}_{2}^{+})$). Note that we do not explicitly verify whether ${\textnormal{z}}_{1}$ itself is correct or incorrect; instead, we infer its correctness from the associated ${\textnormal{a}}_{1}$, which we find to be a reasonable approximation at scale. For online RL (GRPO), we only keep the generated visual problems and its verifiable answer.
Table 2: Main results on vision-centric reasoning benchmarks. We compare our models against both open- and closed-source VLMs across five challenging benchmarks. Community baselines trained on open source data sometimes underperform the base Qwen2.5-VL-7B-Instruct, underscoring the lack of high-quality open reasoning data for vision-centric tasks. Closed source performance is shown for reference, taken from the authors technical reports.
@@
Table 3: Quantitative comparison of Data Complexity and CoT Richness. We compare our Stage 1 and Stage 2 data against the LPT baseline. Lower Pass/Solve rates indicate higher complexity; higher Behavior counts indicate richer reasoning traces.
Method Pass Rate $\downarrow$ Perfect Solve $\downarrow$ CoT Richness $\uparrow$
(Avg. over 8) (8/8 Correct) (Behaviors/Trace)
LPT (Baseline) 66.1% 36.7% 0.65
Ours (Stage 1) 56.8% 25.5% 0.80
Ours (Stage 2) 38.4% 3.3% 1.99
@@ 3 Experiments
In this section, we first compare our synthesized dataset vs a SoTA baseline (LPT) in terms of scale, downstream performance, complexity of the synthesized visual problems, and richness of the distilled CoT trace. Later, we evaluate the performance of finetuning a 7B VLM on our data, and compare it compare against three model/data categories: (i) open-weight, open-data models, (ii) open-weight, closed-data models, and (iii) fully closed models. Next, we use our data to systematically analyze the different post-training stages in a base instruction model, revealing consistent findings. Finally, we evaluate whether our vision-centric reasoning data transfers to a different modality and architecture (e.g. text-only reasoning, embodied QA, and audio reasoning on an Omni-7B model). We also show that the model trained on Long Grounded Thoughts generalizes to open-ended visual problems.
3.1 Scale, Complexity and Richness of CoT
[…]
3.2 Setup and Main Results
@@
Base Model and Benchmarks We fine-tune Qwen2.5-VL-7B-Instruct (Bai et al., 2025) and evaluate on vision-centric tasks. Following (Liao et al., 2025a), for general-knowledge datasets, we retain only their vision splits (e.g., MMStar-V (Chen et al., 2024b)). In addition, we adopt vision-centric focused benchmarks V Bench (Wu and Xie, 2024), CV-Bench (averaged both 2D and 3D), MMVP (Tong et al., 2024) and RealWorldQA (xAI, 2024). These benchmarks test visual search, 2D/3D spatial reasoning, fine-grained attribution, and scene understanding.
@@
Evaluation. We use two main protocols for evaluation. For comparison vs existing models and published results we used VLMEvalKit (Duan et al., 2024) with GPT-4o as a judge (Table 2). For ablations, we utilize the evaluation protocol of  (Liao et al., 2025a) which utilizes rule-based matching instead of a judge. For clarity, results are shown with decimal when no LLM as a judge is used.
Source of high-quality image captions and metadata. We use DOCCI (Onoe et al., 2024), a human-annotated dense caption dataset. We defer exploring other caption datasets and automatically generated captions to future work.
@@
Baselines. We compare against three categories: (i) open-weight, open-data models, (ii) open-weight, closed-data models, and (iii) fully closed models. For open-weight, open-data, we include VLAA-Thinker (Chen et al., 2025b), a model trained on $\sim$152K synthetic reasoning traces using both SFT and RL. We also compare with LPT (Liao et al., 2025a) and ReVisual-R1 (Chen et al., 2025c). Additionally, we include Virgo (Du et al., 2025), trained on $\sim$19K math-heavy datapoints. For open-weight, closed-data, we consider MiMo-VL-7B, a SoTA VLM trained with a four-stage recipe totaling $\sim$2.4T tokens across images, videos, and text. For fully closed models, we report results for GPT-4o and Claude 3.7 taken from (Xiaomi et al., 2025).
Training Algorithms.
@@
SFT. We perform SFT on the large pool of data from Stage 1 and Stage 2. To simplify training, we adopt a curriculum: first fine-tuning on simpler Stage 1 data for one epoch, then introducing the more complex Stage 2 data. Stage 1 consists of $\sim$750K datapoints, and Stage 2 adds $\sim$251K. We fine-tune the language decoder with a batch size of $256$, learning rate $8\times 10^{-6}$. For Stage 1, we train for up to one epoch with maximum image resolution $512\times 512$ and input cutoff length $1024$. For Stage 2, we halve the learning rate and apply early stopping based on validation loss. All experiments are implemented with llama-factory.
Offline RL (DPO). We fine-tune the language decoder with a batch size of $256$ using a learning rates of $\{6\times 10^{-6}\}$. Training runs for up to six epochs with early stopping based on validation loss. For DPO, we set $\beta=1$ and, following (Pang et al., 2024), include the SFT loss with a weight of $0.5$. Implemented with llama-factory.
Online RL (GRPO). We perform GRPO on top of the fine-tuned model via VERL (Sheng et al., 2024). We set the batch size 128, max response length 8192 and the learning rate $1\times 10^{-6}$. We use the KL loss coefficient of $0.001$ and omit entropy penalty. To avoid overfitting to a specific answer format, our labels are set in a diverse format (e.g., (A), A, and (A) answer); thus, we compute the reward based on an LLM judge, assigning reward = $1$ to a correct answer and reward = 0 to a wrong answer (as identified by the judge). We additionally assign format reward, we add $0.1$ to the reward if the response is formatted correct (i.e., the model correctly generates both <think> and </think>).
@@
Main Result. Comparison vs. Baselines. Table 2 compares Qwen2.5-VL-7B fine-tuned with our data against community baselines, close data, and closed models. Models trained on limited open reasoning data (e.g., VLAA-Thinker, Revisual) often underperform even the base Qwen2.5-VL-7B-Instruct, underscoring the lack of high-quality open data for vision-centric reasoning. In contrast, our models consistently improve the base, achieving the best results and outperforming on 3/5 benchmarks.
Generalization to open-ended non-MCQ tasks. Despite leveraging entirely multiple-choice questions, our models generalize robustly to open-ended visual problems and unseen domains. Table 4 shows we achieve a +8.8 point gain on NiEH (Kim and Ammanabrolu, 2025), an open-ended embodied benchmark. We also observe consistent improvements on MathVision (mixed format also including open-ended tasks) (see Table 9). This confirms that our reasoning gains transfer beyond the training format and domain.
Figure 5: Quantitative and qualitative comparison of the post-training pipeline on our data vs pure RL on the base model. (Upper Right) The graph illustrates the effect of scaling dataset size during online RL showcasing relative peak improvement in percentage. The baseline (blue line), starting from an off-the-shelf model, exhibits negative scaling: performance peaks at 0.695 (10K samples) and degrades with more data. In contrast, our method (green line), which includes SFT on our high-quality data with complex reasoning traces, allows to scale online RL further. This suggests that without offline "teaching" of reasoning patterns via SFT, online RL fails to effectively utilize larger datasets. (Upper Left) A qualitative example (from V* bench), using each model’s best checkpoint (indicated by a dot on the curve), highlights the resulting difference in reasoning. The baseline model fails to identify the partially obscured dog and answers incorrectly. Our model also initially expresses confusion but then self-corrects ("Wait, I’m getting conflicting information…") , showcasing a multi-step reasoning process to arrive at the correct answer. This self-correction capability, instilled with our data, is not observed in the baseline, indicating RL alone was insufficient to elicit this behavior. Image brightness was increased for illustration purposes.
@@ 3.3 Analysis of VLM Post-training Stages
We analyze VLM post-training strategies using our data. Table 11 reports aggregated performance, averaged over four benchmarks. Figure 2 shows scaling compared to LPT. For evaluation, we use rule-based matching instead of LLM-as-judge for cost efficiency. Here, we summarize our findings:
Online RL requires prior teaching of cognitive behaviours; starting from the base instruct model underperforms SFT only on our data. When starting from the base model lacking cognitive patterns (see Figures 53(b)), GRPO peaks at 0.695 (10K) and declines with more data (50–100K), remaining below our SFT baseline (0.716). Finetuning first on LPT data improves but still underperforms simple SFT finetuning on our data. Finetuning on our data leads to the best results suggesting that without offline “teaching” of cognitive behaviours, online RL cannot elicit competitive performance vs SFT in high-quality and diverse data. Notably, we do not observed online RL was able to elicit structured complex reasoning traces unless it was pretrained first on our reasoning SFT data (see Figure 5 for a qualitative example).
@@
SFT data diversity matters; grounded MCQ SFT >> LPT SFT. Our grounded-MCQ SFT reaches 0.716 versus 0.682 for LPT SFT (both at comparable scale), a gain of +3.4 points on average.
@@ 3.4 Beyond Vision-Centric Reasoning
Table 4: Out-of-domain evaluation. We assess cross-domain transfer on MMLU-Pro (text-only) and NiEH (embodied QA). While baselines like Virgo and VLAA degrade text reasoning (negative transfer), our method preserves performance. Additionally, our model generalizes effectively to the open-ended single evidence NiEH benchmark, achieving a +8.8 point gain over the base model despite using only MCQ training data.
Model MMLU-Pro NiEH EQA
Qwen2.5-VL 47.15 47.55
VLAA-thinking 21.56 47.85
@@
LPT 50.77 51.95
Ours (SFT) 50.82 48.24
Ours (SFT+GRPO) 47.00 39.10
Ours (SFT+DPO) 47.07 56.34
Table 5: Omni-modality out-of-distribution results on audio MMAU Benchmark(Sound, Music, Speech) and text-only (MMLU-Pro). We use Qwen2.5-Omni-7B as baseline.
@@
Table 4 shows our model improves upon the base model and existing baselines on text-only reasoning MMLU-PRO (Wang et al., 2024), proving our data does not harm text-only capabilities. Further, despite no video/embodied training, we achieve $\sim$ 8-point gains on the NiEH single-evidence. Notably, this is also an open-ended VQA tasks (Kim and Ammanabrolu, 2025). Setup details are in Appendix A.6.3. Table 5, on the other hand, shows that fine-tuning a Qwen2.5-Omni-7B on our data yields consistent cross-modal gains: +0.98 (Sound), +3.02 (Music), and +0.33 (Speech) on MMAU (Sakshi et al., 2024), plus +4.07 on MMLU-Pro. The improvement on MMLU-Pro is more notable in the Omni-model than in the base Qwen2.5-VL. These results demonstrate robust positive transfer from vision-centric data to audio and text modalities.
@@ 4 Related Work
Recent efforts have been made to enhance visual reasoning in VLMs, spanning from test-time techniques (Liao et al., 2024, 2025b; Acuna et al., 2025) to improved architectural designs and training pipelines (Cheng et al., 2024; Wu et al., 2025; Chen et al., 2025a). However, most efforts to distill high-quality reasoning datasets at scale remain largely in the text-only domain (Team, 2025; Jung et al., 2025; Lu et al., 2026; Kim et al., 2026). In contrast, efforts to build multimodal datasets remain largely linear and/or utilize existing questions rather than synthesizing new problems with different levels of complexity covering the entire post-training VLM pipeline at scale (Xu et al., 2024; Ma et al., 2025; Corbière et al., 2025; Ishaq et al., 2025). A smaller body of work has begun studying more complex reasoning structures such as reflection, self-correction, or iterative refinement in VLMs. For example, Virgo (Du et al., 2025) distills multimodal CoTs from multimodal reasoning models (QwenTeam, 2024), focusing on math-heavy domains. VLAA-Thinking (Chen et al., 2025b) similarly distills CoTs in the general domain, though subsequent SFT on these traces has been shown to degrade performance. LPT (Liao et al., 2025a) introduced long-form CoTs distilled from large reasoning models, such as R1 (DeepSeek-AI et al., 2025), with explicit emphasis on expanding thoughts and encouraging models to deliberate more thoroughly across perception-heavy tasks. Our work builds on this line. However, rather than restricting CoTs to fixed stages, linear narratives or reusing existing visual questions, we synthesize diverse and verifiable visual problems with different levels of complexity and complex reasoning traces, scaling along three axes: scale, complexity, and reasoning depth.